Find the Hamiltonian and write the canonical equations of a harmonic oscillator, if its Lagrangian is L=(x ̇ 2 – ω2x2)




1
Expert's answer
2021-06-18T11:23:41-0400

Hamiltonian in terms of Lagrangian is "H = p \cdot x - L", where "p = \frac{\partial L}{\partial{\dot x}} = 2 \dot{x}". "H" is a function of "p, x".

Hence, "H = 2 \dot{x} \cdot \dot{x} - \dot{x}^2 + \omega^2 x^2 = \dot{x}^2 + \omega^2 x^2 = \frac{p^2}{4} + \omega^2 x^2".

Hamiltonian equations are "\dot{x} = \frac{\partial H}{\partial p}", "\dot{p} = - \frac{\partial H}{\partial x}".

Using Hamilton above: "\dot{x} = \frac{p}{2}", "\dot{p} = -2 \omega^2 x".

Substituting "p = 2 \dot{x}" from the first equation into the second, obtain "2 \ddot{x} = - 2 \omega^2 x", or "\ddot{x} + \omega^2 x = 0", which is the equation of motion of harmonic oscillator.